# Reinforcing the Diffusion Chain of Lateral Thought with Diffusion Language Models

### v2

@@ 1 Introduction
We study two paradigms of diffusion language models to reinforce the DCoLT, continuous-time diffusion language models [24, 6] and discrete-time diffusion language models [3, 19, 27, 42]. For the continuous-time paradigm, we consider SEDD [24] as a representative DLM. SEDD predicts the concrete score, allowing for a closed-form expression of the predicted diffused distribution of generated tokens at each step. This distribution can be viewed as a probabilistic policy for sampling tokens, which can be trained to optimize the reward on the final answers by reinforcing DCoLT.
@@
We conduct experiments on both Math and code generation tasks to demonstrate the efficacy of DCoLT. After training the DCoLT on the SEDD 400M model, it achieves 96.2% and 57.0% in accuracy on Sudoku 4$\times$4 and GSM8K-Aug [10] tasks. On the LLaDA 8B model, DCoLT achieves the state-of-the-art performance among existing DLMs trained with SFT or RL or even both. Using only public data and 16 H800 GPUs, the DCoLT-reinforced LLaDA model achieves 88.1% on GSM8K [8], 44.6% on MATH [16], 51.6% on MBPP [4] and 59.1% on HumanEval [7] in the challenging zero-shot setting. Even compared with auto-regressive models that are trained with significantly more proprietary data and fully annotated CoT reasoning processes, it still demonstrates competitive performances.
@@ 2 Related Works
Diffusion models have achieved impressive results in image modeling [18, 37, 33, 21, 43], with recent efforts [24, 11, 13, 3, 48, 2, 47, 12, 29, 27, 42] extending to language tasks. Depending on the types of diffused distributions over text tokens, diffusion language models can be categorized into continuous diffusion models [13] and discrete diffusion models [3, 24, 11], with our work focusing primarily on the latter due to the discrete nature of language.
Among discrete diffusion models, masked diffusion models [27, 42, 24] emerge as a promising approach due to their superior performances. Recent advances simplify model training and design by adopting the straight cross-entropy loss [48, 29] and removing the time embeddings [47]. The resultant models [27, 42] achieve the state-of-the-art performance among existing diffusion language models.
@@ 3.1 Diffusion Chain of Lateral Thought
6:         for $n=1$ to $N$ do $\triangleright$ $N$ denotes the number of denoising steps
@@
10:                 Calculcate the ranking score $h_{\theta,n}$ for each token
@@
15:         $r^{g}=\operatorname{reward\_func}(q,x^{g}_{N})$triangleright Compute the rewards
@@
20:    for $n=1$ to $N$ do $\triangleright$ Compute $\pi_{\theta}$ and losses for each denoising step
@@
22:             $\pi_{\theta,n}(x^{g}_{n}|x^{g}_{n-1})=\prod_{i=1}^{|x^{g}_{n}|}p_{\theta,t_{n}}(x_{n}^{g,i}|x^{g}_{{n-1}})$triangleright see Eq. 5
@@
24:             $\pi_{\theta,n}(x^{g}_{n}|x^{g}_{n-1})=\pi_{\theta,n}^{\text{unmask}}(\mathcal{U}^{g}_{n}|x_{n})\cdot\pi_{\theta,n}^{\text{token}}(x_{n}|x_{n-1},\mathcal{U}_{n})$triangleright see Eq. 9
@@ 3.2 A Continuous-Time DLM Case: DCoLT-reinforced SEDD
$p_{\theta,t_{n}}(x_{n}^{i}|x_{n-1})$ =$=\begin{cases}s_{\theta}(x_{{n-1}},t_{n-1})_{i,x_{n}^{i}}\cdot(t_{n-1}-t_{n})\cdot Q_{t_{n-1}}(x_{n-1}^{i},x_{n}^{i}),&x_{n-1}^{i}\neq x_{n}^{i},\\ 1-\sum_{y\neq x_{n}^{i}}s_{\theta}(x_{{n-1}},t_{n-1})_{i,y}\cdot(t_{n-1}-t_{n})\cdot Q_{t_{n-1}}(x_{n-1}^{i},y),&x_{n-1}^{i}=x_{n}^{i},\end{cases}$ (6)
where $|x_{n}|$ denotes the token length of the sequence $x_{n}$. Note that the predicted $p_{t_{n}}$ is dependent on the model parameters $\theta$. During training, we may update $\theta$ to reinforce or suppress the probability of all actions along a trajectory of thoughts $[x_{0},x_{1},...,x_{N}]$, according to the reward associated with each completion. Unlike traditional diffusion model training, which typically optimizes each individual denoising step independently with ground truth responses, DCoLT jointly optimizes the entire reverse process of thought from $t_{0}=T$ to $t_{N}=0$, promoting the emergence of lateral thought leading to correct answers.
@@ 4.1.2 Results
Particularly, SEED + DCoLT achieves an accuracy of 96.2% on Sudoku $4\times 4$ task, the best performance among all the methods. Using the same SEED model, DCoLT greatly exceeds DoT ($79.4\%$), even though the latter introduce step-by-step annotated CoT data for supervised training. For GPT2 model, both CoT and Implicit CoT, no matter if being post-trained with SFT or RL, perform worse than our model clearly. On the GSM8K-Aug dataset, SEED + DCoLT also reaches $57.0\%$ in accuracy, which is notably higher than both DoT and CoT counterparts. We will further analyze the generation process of SEED + DCoLT in Appendix A, demonstrating how tokens are generated in a lateral thinking manner.
@@
Models Post-Training Sudoku 4$\times$4 GSM8K-Aug
@@ 4.2.1 Settings
We further apply DCoLT to LLaDA [27], a SOTA discrete-time masked-based DLM with 8B parameters. During generation, it starts with a masked sequence of length $256$, and performs a 256-step reverse diffusion process. We adopt a block-wise unmasking strategy as in LLaDA [27], and divide the sequence into blocks of length $8$. In Tab. 2, both LLaDA and LLaDOU (LLaDA+DCoLT) are evaluated with this setting for fair comparison. We conduct experiments with 16 H800 GPUs to jointly train both UPM and LLaDA parts. For implementation details, please refer to Appendix C.2.
@@ 4.2.2 Results
Even when compared with auto-regressive models trained with a significantly larger amount of proprietary data, LLaDOU remains highly competitive. For example, DeepseekMath [36] gathers 776K questions with reasoning process annotations for SFT and 144K questions for RL, while LLaDOU is trained on just 15K public training samples from GSM8K and MATH. Despite using nearly two orders of magnitude fewer data for RL training, LLaDOU matches DeepseekMath’s performance on GSM8K (88.1%), highlighting its data efficiency and the effectiveness of reward-driven reasoning via diffusion-based lateral reasoning. In Appendix B, we analyze the thinking process generated by LLaDOU.
Similar trends are observed on code generation benchmarks. Unlike models that rely on valuable ground-truth code for supervised training, LLaDOU requires no code for supervised training at all. Instead, it only needs several test cases to provide a simple outcome-based reward for reinforcement learning: the model is rewarded only if the generated code passes all unit tests. Despite this code-free reward signal, LLaDOU achieves 51.6% on MBPP and 59.1% on HumanEval, outperforming other DLMs and auto-regressive models, except for DeepseekCoder [14] that benefits from 2B tokens of high-quality instruction-tuning data—approximately 150×150$\times$150 × more than we used (about 12M tokens). Details of post-training datasets used by all compared methods are provided in Appendix C.3.
@@ 4.2.3 Ablation Studies
We provide some ablation studies to reveal the role of some model design and hyper-parameter settings in LLaDOU. Unless otherwise specified, all experiments are conducted on LLaDOU 8B with $N=64$ diffusion steps, and the model is trained for $150$ iterations with a batch size of $32$.
@@
Though the models in Tab. 2 is trained to generate completions with a fixed generation length $256$ (i.e., the length of initialized mask tokens), it can benefit from generating longer sequences without further RL training on longer generations. This is especially effective when answering difficult questions on MATH dataset, in which solving problems sometimes require longer reasoning. Tab. 4 reports the accuracies with different generation lengths. It suggests that longer generations in inference lead to better accuracies. For example, a generation length of $384$ achieves an accuracy of 46.22%, which is +1.62% higher than the default length of $256$. This improvement vanishes as we further increase the generation length. However, if we further tune the model on a longer generation length, for example $L=512$, the accuracy further improves to 47.3%. This suggests the potential of how scaled generation lengths may improve the performance of DCoLT-trained DLMs, as shown in Fig. 13 of Appendix D.
@@
LLaDA + DCoLT (LLaDOU) 8B 44.6 45.7 46.2 45.7 45.9
@@ C.2.1 Implementation Details for Math Problems
We assign a hard reward for each completion, that it is $1$ only if the final answer is equivalent to the ground truth [16].
@@
The model is trained with 64 prompts in a batch, each generating 16 completions to form a group for advantage calculation. We take an AdamW optimizer with a learning rate of $5\times 10^{-6}$, and $(\beta_{1},\beta_{2})=(0.9,0.999)$. We do not apply the KL penalty by default, as it provides marginal benefits in our experiments. The whole training lasts for $140$ iterations on 16 H800 GPUs, which takes about 63 GPU days (i.e., about 4 days on wall clock with 16 GPUs).
@@ Appendix D Potential Scaling Law for Longer Generations
(a) The graphs of $y=x^{4}$ and $y=7x^{2}-10$ intersect at four points with $x$-coordinates ±mplus-or-$\pm\sqrt{m}$ and ±$\pm\sqrt{n}$, where $m>n$. What is $m-n$?
(a) Find $4321_{5}-1234_{5}$. Express your answer in base $5$.

### v3

@@ Abstract
footnotetext: This project was initiated and supported by MAPLE Lab at Westlake University.
@@ 1 Introduction
We study two paradigms of diffusion language models to reinforce the DCoLT, continuous-time diffusion language models [24, 6] and discrete-time diffusion language models [3, 19, 27, 43, 37]. For the continuous-time paradigm, we consider SEDD [24] as a representative DLM. SEDD predicts the concrete score, allowing for a closed-form expression of the predicted diffused distribution of generated tokens at each step. This distribution can be viewed as a probabilistic policy for sampling tokens, which can be trained to optimize the reward on the final answers by reinforcing DCoLT.
@@
We conduct experiments on both Math and code generation tasks to demonstrate the efficacy of DCoLT. After training the DCoLT on the SEDD 400M model, it achieves 96.2% and 57.0% in accuracy on Sudoku $\times4$ and GSM8K-Aug [10] tasks. On the LLaDA 8B model, DCoLT achieves the state-of-the-art performance among existing DLMs trained with SFT or RL or even both. Using only public data and 16 H800 GPUs, the DCoLT-reinforced LLaDA model achieves 88.1% on GSM8K [8], 44.6% on MATH [16], 51.6% on MBPP [4] and 59.1% on HumanEval [7] in the challenging zero-shot setting. Even compared with auto-regressive models that are trained with significantly more proprietary data and fully annotated CoT reasoning processes, it still demonstrates competitive performances.
@@ 2 Related Works
Diffusion models have achieved impressive results in image modeling [18, 38, 33, 21, 44], with recent efforts [24, 11, 13, 3, 49, 2, 48, 12, 29, 27, 43, 37] extending to language tasks. Depending on the types of diffused distributions over text tokens, diffusion language models can be categorized into continuous diffusion models [13] and discrete diffusion models [3, 24, 11], with our work focusing primarily on the latter due to the discrete nature of language.
Among discrete diffusion models, masked diffusion models [27, 43, 24] emerge as a promising approach due to their superior performances. Recent advances simplify model training and design by adopting the straight cross-entropy loss [49, 29, 37] and removing the time embeddings [48]. The resultant models [27, 43] achieve the state-of-the-art performance among existing diffusion language models.
@@ 3.1 Diffusion Chain of Lateral Thought
6:    for $n=1$ to NN do $\triangleright$ NN denotes the number of denoising steps
@@
10:      Calculate the ranking score $h_{\theta,n}$ for each token
@@
15:    $r^{g}=\operatorname{reward\_func}(q,x^{g}_{N})$ $\triangleright$ Compute the rewards
@@
20:  for $n=1$ to NN do $\triangleright$ Compute $\pi_{\theta}$ and losses for each denoising step
@@
22:     $\pi_{\theta,n}(x^{g}_{n}|x^{g}_{n-1})=\prod_{i=1}^{|x^{g}_{n}|}p_{\theta,t_{n}}(x_{n}^{g,i}|x^{g}_{{n-1}})$ $\triangleright$ see Eq. 5
@@
24:     $\pi_{\theta,n}(x^{g}_{n}|x^{g}_{n-1})=\pi_{\theta,n}^{\text{unmask}}(\mathcal{U}^{g}_{n}|x_{n})\cdot\pi_{\theta,n}^{\text{token}}(x_{n}|x_{n-1},\mathcal{U}_{n})$ $\triangleright$ see Eq. 9
@@ 3.2 A Continuous-Time DLM Case: DCoLT-reinforced SEDD
$p_{\theta,t_{n}}(x_{n}^{i}|x_{n-1})$ $=\begin{cases}s_{\theta}(x_{{n-1}},t_{n-1})_{i,x_{n}^{i}}\cdot(t_{n-1}-t_{n})\cdot Q_{t_{n-1}}(x_{n-1}^{i},x_{n}^{i}),&x_{n-1}^{i}\neq x_{n}^{i},\\ 1-\sum_{y\neq x_{n}^{i}}s_{\theta}(x_{{n-1}},t_{n-1})_{i,y}\cdot(t_{n-1}-t_{n})\cdot Q_{t_{n-1}}(x_{n-1}^{i},y),&x_{n-1}^{i}=x_{n}^{i},\end{cases}$ (6)
where $|x_{n}|$ denotes the token length of the sequence $x_{n}$. Note that the predicted $p_{t_{n}}$ is dependent on the model parameters $\theta$. During training, we may update $\theta$ to reinforce or suppress the probability of all actions along a trajectory of thoughts $x_{0},x_{1},...,x_{N}]$, according to the reward associated with each completion. Unlike traditional diffusion model training, which typically optimizes each individual denoising step independently with ground truth responses, DCoLT jointly optimizes the entire reverse process of thought from $t_{0}=T$ to $t_{N}=0$, promoting the emergence of lateral thought leading to correct answers.
@@ 4.1.2 Results
Particularly, SEDD + DCoLT achieves an accuracy of 96.2% on Sudoku $4\times 4$ task, the best performance among all the methods. Using the same SEDD model, DCoLT greatly exceeds DoT ($79.4\%$), even though the latter introduce step-by-step annotated CoT data for supervised training. For GPT2 model, both CoT and Implicit CoT, no matter if being post-trained with SFT or RL, perform worse than our model clearly. On the GSM8K-Aug dataset, SEDD + DCoLT also reaches $57.0\%$ in accuracy, which is notably higher than both DoT and CoT counterparts. We will further analyze the generation process of SEDD + DCoLT in Appendix A, demonstrating how tokens are generated in a lateral thinking manner.
@@
Models Post-Training Sudoku $\times4$ GSM8K-Aug
@@ 4.2.1 Settings
We further apply DCoLT to LLaDA [27], a SOTA discrete-time masked-based DLM with 8B parameters. During generation, it starts with a masked sequence of length $256$, and performs a 256-step reverse diffusion process. We adopt a block-wise unmasking strategy as in LLaDA [27], and divide the sequence into blocks of length $8$. In Tab. 2, both LLaDA and LLaDOU are evaluated with this setting for fair comparison. We conduct experiments with 16 H800 GPUs to jointly train both UPM and LLaDA parts. For implementation details, please refer to Appendix C.2.
@@ 4.2.2 Results
Even when compared with auto-regressive models trained with a significantly larger amount of proprietary data, LLaDOU remains highly competitive. For example, DeepseekMath [36] gathers 776K questions with reasoning process annotations for SFT and 144K questions for RL, while LLaDOU is trained on just 15K public training samples from GSM8K and MATH. Despite using nearly two orders of magnitude fewer data for RL training, LLaDOU matches DeepseekMath’s performance on GSM8K (88.1%), highlighting its data efficiency and the effectiveness of reward-driven reasoning via diffusion-based lateral reasoning. In Appendix B, we provide a detailed analysis of the thinking process generated by LLaDOU, including both visualization examples and quantitative evaluations. Fig. 12 illustrates the reward curve during DCoLT training in the math domain, showing an increasing trend in the outcome-based reward over training iterations.
Similar trends are observed in code generation benchmarks. Unlike models that rely on valuable ground-truth code for supervised training, LLaDOU requires no code for supervised training at all. Instead, it only needs several test cases to provide a simple outcome-based reward for reinforcement learning: The model is rewarded only if the generated code passes all unit tests. Despite this code-free reward signal, LLaDOU achieves 51.6% on MBPP and 59.1% on HumanEval, outperforming other DLMs and auto-regressive models, except for DeepseekCoder [14] that benefits from 2B tokens of high-quality instruction-tuning data—approximately $150\times$ more than we used (about 12M tokens). Details of post-training datasets used by all compared methods are provided in the Appendix C.3.
Table 3: Accuracy of LLaDOU on MATH subsets across difficulty levels (5: the hardest and 1: the easiest) and generation lengths. The result at $L=512$ is obtained from the model fine-tuned with this length, while results at other lengths are directly evaluated using LLaDOU trained with $L=256$.
Length Level 1 Level 2 Level 3 Level 4 Level 5
128 80.8% 61.0% 45.6% 29.5% 13.2%
256 83.3% 65.1% 52.3% 35.9% 18.7%
384 82.4% 66.6% 54.9% 39.0% 20.2%
512 82.6% 69.7% 56.9% 40.2% 21.5%
@@ 4.2.3 Ablation Studies
We provide some ablation studies to reveal the role of some model design and hyper-parameter settings in LLaDOU. Unless otherwise specified, all experiments are conducted on LLaDOU with $N=64$ diffusion steps, and the model is trained for $150$ iterations with a batch size of $32$.
@@
Though the models in Tab. 2 is trained to generate completions with a fixed generation length $256$ (i.e., the length of initialized mask tokens), it can benefit from generating longer sequences without further RL training on longer generations. This is especially effective when answering difficult questions on MATH dataset, in which solving problems sometimes require longer reasoning. Tab. 5 further summarizes the overall results under different generation lengths and models. It suggests that longer generations in inference lead to better accuracies. For example, a generation length of $384$ achieves an accuracy of 46.22%, which is +1.62% higher than the default length of $256$. This improvement vanishes as we further increase the generation length. However, if we further tune the model on a longer generation length, for example $L=512$, the accuracy further improves to 47.3%.
In Tab. 3, we report the accuracy of LLaDOU on MATH subsets across difficulty levels. The accuracy on level-1 problems saturates with length 256, suggesting that shorter generations are sufficient for simpler questions. In contrast, the performance on harder problems (levels 2–5) continues to improve with longer generations, indicating that complex reasoning benefits from longer responses. These results suggest the potential of how scaled generation lengths may improve the performance of DCoLT-trained DLMs, as shown in Fig. 14 of Appendix D.
@@
LLaDOU (LLaDA + DCoLT) 8B 44.6 45.7 46.2 45.7 45.9
@@ 5 Conclusion
Acknowledgements
This work was supported by National Natural Science Foundation of China under Grant No. 92467104, and Zhejiang Leading Innovative and Entrepreneur Team Introduction Program (2024R01007).
@@ C.1 SEDD+DCoLT
For the Sudoku $4\times 4$ task, the reward is determined by verifying whether the generated solution satisfies all the rules of a valid $4\times 4$ Sudoku, namely that each row, each column, and each $2\times 2$ subgrid must contain the digits 1–4 exactly once. A reward of 1 is assigned only when the solution is entirely correct; otherwise, the reward is 0.
@@ C.2.1 Implementation Details for Math Problems
[…]
@@
The model is trained with 64 prompts in a batch, each generating 16 completions to form a group for advantage calculation. We take an AdamW optimizer with a learning rate of $5\times 10^{-6}$, and $(\beta_{1},\beta_{2})=(0.9,0.999)$. We do not apply the KL penalty by default, as it provides marginal benefits in our experiments. The whole training lasts for $140$ iterations on 16 H800 GPUs, which takes about 63 GPU days (i.e., about 4 days on wall clock with 16 GPUs). We visualize the reward curves during training in Fig. 12.
Figure 12: Reward curves on MATH and GSM8K during DCoLT training. Dashed lines denote the raw rewards, while solid lines represent the moving-average smoothed rewards, showing the overall trend.
@@ C.2.2 Implementation Details for Code Generation
C.2.3 Direct Comparison with d1-LLaDA
A direct comparison with d1[47] using the same LoRA structure and dataset still reveals the superior performance by our method. Specifically, we train LLaDOU on GSM8K only, using LoRA with rank $r=128$ and scaling factor $\alpha=64$. During inference, we set the generation length to 256, the number of diffusion steps to 128, and the block length to 32 — mirroring d1’s configuration. In the same setting, DCoLT with LoRA achieves an accuracy of 84.7%, which outperforms d1 with both diffu-GRPO (79.8%) and d1-LLaDA (81.1%) setups, as shown in Tab. 8. These results indicate that LLaDOU still performs better than d1 with the same LoRA structure and the training data.
Table 8: Comparison between d1 and LLaDOU with LoRA on GSM8K. Both methods only use the GSM8K to train the LLaDA model with their respective RL approach.
Model Accuracy (%)
LLaDA 8B 78.3
   +diffu-GRPO 79.8
   d1-LLaDA (SFT + diffu-GRPO) 81.1
LLaDOU 8B (LoRA) 84.7
C.2.4 Applying DCoLT to Dream-7B on GSM8K
To further examine the generality of DCoLT, we apply it to the Dream-7B [43] model on GSM8K. Unlike LLaDA, which is a natively masked diffusion model, Dream-7B originates from an auto-regressive model.
Using the same reinforcement learning configuration as in the main experiments (64 denoising steps and generation length $L=256$), DCoLT raises Dream-7B’s accuracy from 50.11% to 80.53%, an absolute gain of +30.42%. The improvement magnitude is comparable to that observed on LLaDA-8B, suggesting that DCoLT consistently enhances reasoning ability across heterogeneous diffusion architectures. The model is trained with a batch of 64 prompts, each generating 16 completions for advantage calculation, using AdamW with a learning rate of $2.5\times 10^{-6}$ and $(\beta_{1},\beta_{2})=(0.9,0.999)$.
@@ Appendix D Potential Scaling Law for Longer Generations
(a) The graphs of $y=x^{4}$ and $y=7x^{2}-10$ intersect at four points with $x$-coordinates $\pm\sqrt{m}$ and $\pm\sqrt{n}$, where $m>n$. What is mnm-n?
(a) Find $4321_{5}-1234_{5}$. Express your answer in base 55.